OpenStudy (anonymous):

A spider has 8 legs. An ant has 6 legs. There are a group of spiders and a group of ants. The grups has equal numbers of legs. What is the least number of spider and ants?

OpenStudy (anonymous):

3 spiders and 4 ants

OpenStudy (anonymous):

3*8=24 4+6=24 so 3 spiders, and 4 ants

OpenStudy (turingtest):

8x=6y where x=# spiders and y=# ants 6/8=y/x=3/4 y=4, x=3

OpenStudy (anonymous):

It would be 6/8=x/y not y/x.

OpenStudy (anonymous):

I'm very bad at maths. Could you please explain in details how you got the answer

OpenStudy (anonymous):

Ok whatch it like this, one spider=8 legs, and one ant=6 legs u can imagine that x is number of spiders and y is number os ants.. so number of spider legs is 8*x, and number of ant legs is 6*y Question is to find x and y, but u know that number of legs has to be same so: 8*x=6*y, and u just have to find 2 numbers that matches(u dont even have to do any calculating...)